OpenStudy (anonymous):

Charlie is at a small airfield watching for the approach of a small plane with engine trouble. He sees the plane at an angle of elevation of 32 . At the same time, the pilot radios Charlie and reports the plane’s altitude is 1,700 feet. Charlie’s eyes are 5.2 feet from the ground. Draw a sketch of this situation (you do not need to submit the sketch). Find the ground distance from Charlie to the plane. Explain your work.

OpenStudy (hba):

k did u work out the problem

OpenStudy (anonymous):

This what I did I'm i right we see that one vertical hypotenuse a is1,700 feet and so we have to solve horizontal hypotenuse b. We can use tangent function tan(32) = a/b => b = a/tan(32) ~ 2570 feet.

OpenStudy (anonymous):

whats your explanation.

OpenStudy (hba):

vertical is 1700 feet and unknown angle is 32 degree tan 32=1698.4/H subracting height from altitude produces a right triangle H=2712 feet away

OpenStudy (anonymous):

I think yours is better lol

OpenStudy (hba):

i think u have 15 seconds to get out of way

OpenStudy (anonymous):

okay ... thanks for your input

OpenStudy (hba):

ur welcome